OpenStudy (happinessbreaksbones):

factoring 2x^2 - 7x - 4 = 0 ?? for some reason I'm stuck on this one.. I need to find the solutions.

OpenStudy (lgbasallote):

\[2x^2 - 7x - 4 = 0\] break down the middle term \[(2x-8x + x - 4 = 0\] factor out \[2x(x - 4) + (x-4) = 0\] factor out again \[(x-4)(2x+1) = 0\] does that help?

OpenStudy (happinessbreaksbones):

yeah the final answer (x - 4)(2x +1) was what I had gotten.. but I need to find the solutions for x after that :/ I know x - 4 = 0 +4 +4 x= 4 but I'm trying to figure out what x equals for (2x + 1) thank you

OpenStudy (lgbasallote):

\[2x + 1 = 0\] \[2x = -1\] \[x = -\frac 12\]

OpenStudy (lgbasallote):

got it?

OpenStudy (happinessbreaksbones):

ohh so it turns into a fraction.. Okay thank you :)

OpenStudy (lgbasallote):

<tips imaginary hat>